Solve the following systems of equations:
7x-6y= -11
9x-2y= 3
Also, can you please explain how to do a problem like this? That'd be great if you did! Thanks, in advance.

Answer by nerdybill(7384)

starting with:
7x-6y= -11
9x-2y= 3
.
we can apply the "addition method" by multiplying the second equation by -3:
7x-6y= -11
-27x+6y= -9
.
Now, we add the two equations together:
7x-6y= -11
-27x+6y= -9
-------------
-20x = -20
x = 1
.
substitute the above into either of the two equations to find y:
9x-2y= 3
9(1)-2y= 3
9-2y= 3
-2y= -6
y = 3
.
Solution: (x,y) = (1,3)